# pAGN: the one-stop solution for AGN disc modeling

### v1

@@ 1 Introduction
[…]
@@
AGN discs are unique astrophysical environments with a rich phenomenology, including high-energy jets, dusty torii, and accreting BHs (Padovani et al., 2017). In the context of gravitational-wave observations, AGN discs are studied as host environments for compact-binary formation and mergers (McKernan et al., 2012, 2011; Yang et al., 2019; Secunda et al., 2019; Fabj et al., 2020; Tagawa et al., 2020; Trani et al., 2023). The large escape velocity around a supermassive BH implies that objects are likely to be retained in the disc vicinities, potentially forming a large population of stellar-mass BHs that have a higher likelihood of interacting. The dense gas in the disc can facilitate binary formation, accelerate the inspiral, and induce chains of hierarchical BH mergers (Gerosa & Fishbach, 2021; Santini et al., 2023; Whitehead et al., 2023; Vaccaro et al., 2023). The occurrence of hierarchical mergers in AGN discs crucially depends on the presence of the so-called migration traps, namely locations in the disc where the migration torque changes sign, which is still an open issue in AGN-disc modeling (Bellovary et al., 2016; Tagawa et al., 2020; Grishin et al., 2023).
@@
Early models of AGNs discs consists of one-dimensional, steady-state, semi-analytic solutions utilizing parametric prescriptions. Subsequent computational advancements allowed for models capturing more complex physics, such as radiative transfer, gas phase transitions, magnetic fields, and general relativity (e.g. Wilson, 1972; Hawley et al., 1984; Falle, 1991; Schartmann et al., 2005; Schartmann et al., 2008; Wada et al., 2009; Huško & Lacey, 2023). Nevertheless, one-dimensional models remain highly valuable today due to their computational efficiency and insightful perspectives on the structure of AGN discs. This makes them particularly useful in the study of interactions between compact objects and BHs. The first of these one-dimensional approaches dates back to Shakura & Sunyaev (1973), who first model geometrically thin, optically thick discs around a BHs. Building on this seminal work, two models are most commonly used in the field, namely those by Sirko & Goodman (2003) and Thompson et al. (2005) (but see also Goodman 2003; Haiman et al. 2009; Cantiello et al. 2021; Gilbaum & Stone 2022; Grishin et al. 2023). Both these models assume some heating mechanisms in the disc that marginally support the outer regions from collapsing due to self-gravity and formulate one-dimensional sets of equations for the AGN-disc profile as a function of a number of parameters such as the mass of the central BH and the accretion rate.
@@
The Sirko & Goodman (2003) and Thompson et al. (2005) models are widely used and underpin some of the key, qualitative results in the field of AGN-disc physics. Despite that, the underlying parameters and methods are often left unspecified. Achieving a stable numerical implementation of these disc solutions is not straightforward and codes in this area have been kept proprietary. The goal of this paper is to critically re-analyze the AGN disc models by Sirko & Goodman (2003) and Thompson et al. (2005). In particular, we clarify the model equations one needs to solve (and crucially the order one needs to solve them), highlight the choices one has to make to obtain stable solutions, and provide a highly customizable implementation. Our software is made publicly available in a Python packaged called pAGN (short for “parametric AGNs”, pronounced as “pagan”).
@@ 2.1.2 Inner regime
The temperature profile in the disc depends on the optical depth $\tau_{\rm v}=\kappa\rho h$, where $\kappa(\rho,T)$ is the opacity. The latter is obtained using interpolated values by Semenov et al. (2003) when $T<10^{4}{\rm K}$ and Badnell et al. (2005) when $T>10^{4}{\rm K}$, the set of which we refer to as the “combined” opacity. These are newer prescriptions for the opacity compared to those by Iglesias & Rogers (1996) and Alexander & Ferguson (1994) used by Sirko & Goodman (2003). From the opacity and effective temperature, we look for solutions in $T$ by assuming the disc ring is in radiative equilibrium:
@@ 2.1.3 Outer regime
Figure 3: Radial profile in the SG03 disc model for the opacity $\kappa$, the mass density $\rho$, the pressure fraction $\beta$, the optical depth $\tau_{\rm v}$, the disc mass $M_{\rm disc}$, the sound speed $c_{\mathrm{s}}$, the temperature $T$, the most unstable mass $M_{\rm{most}\,\rm{unst.}}$, the Toomre stability parameter $Q_{\mathrm{S}}$, the effective temperature $T_{\rm eff}$, the surface mass density $\Sigma_{\mathrm{g}}$ and the half-thickness $h$. All input values are the same as in Fig. 2 by Sirko & Goodman (2003): $M=10^{8}M_{\odot}$, $\alpha=0.1$, $l_{\mathrm{E}}=0.5,\epsilon_{\rm S}=0.1$. Blue (orange) curves indicate the case where $b=0$ ($b=1$) and the viscosity is proportional to total (gas) pressure.
@@ 2.1.4 Disc profiles
[…]
@@ 2.2.4 Accretion criterion
Unlike Sirko & Goodman (2003), the Thompson et al. (2005) model presents an accretion rate $\dot{M}$ that changes as a function of the radial separation $r$. This naturally means that if the accretion rate at the outer boundary is too low, not enough gas is able to reach the central BH to maintain high temperatures and bright AGN luminosities, which are expected to be in the $10^{-5}-10^{-3}L_{\rm Edd}$ range, see Heckman et al. (2004); Kauffmann & Heckman (2009); Trump et al. (2015); Kong & Ho (2018). This introduces a minimum threshold for $\dot{M}_{\rm out}$. Thompson et al. (2005) argue that accretion rates of ${\sim}1-10\rm\,M_{\odot}{\rm yr}^{-1}$ at the inner disc boundary $r_{\rm min}$ are sufficient to produce a bright AGN when the central BH mass is ${\sim}10^{9}\rm\,M_{\odot}$. This is equivalent to a minimum BH accretion rate of $\dot{M}\sim 0.2~{}\dot{M}_{\rm Edd}=0.2~{}L_{\rm Edd}/(0.1c^{2})$ at $r=r_{\rm min}$. Using Eq. (47) in Thompson et al. (2005), we find that over a wavelength range of $[10^{-8}\,\mathrm{m},10^{-3}\,\mathrm{m}]$, setting $0.2~{}\dot{M}_{\rm Edd}$ gives a disc bolometric luminosity of $2\times 10^{-4}\,L_{\rm Edd}$.
@@
We find that using $\dot{M}_{\mathrm{c}}$ from Eq. (26) as a threshold is too stringent and often omits signals that produce bright AGN discs. In the following, we use Eq. (27) as an initial guess for $\dot{M}_{\rm out}$ but then make adjustments if the accretion rate is not large enough to form a luminous AGN. Developing a full prescription is left to future work. As a precaution to avoid setting an $\dot{M}_{\rm out}$ that is too high, TQM05 suggests a maximum limit for $\dot{M}_{\rm out}$ equal to $\dot{M}_{\rm max}=8\pi\rho h\sigma^{2}r/\epsilon_{\rm T}c=L_{\rm M}/\epsilon_{\rm T}c^{2}$, where $L_{\rm M}$ is the Eddington-like luminosity for a galaxy whose momentum driven winds are not high enough to significantly reduce the gas in the disc (Murray et al., 2005).
@@ 2.2.5 Disc profiles
[…]
@@
[…]
@@ 3.2 Input parameters
The SG03 model has five input parameters: the mass of the central BH $M$, the luminosity ratio $l_{E}$ (or alternatively the accretion rate $\dot{M}$), the disc viscosity $\alpha$, the BH radiative efficiency $\epsilon_{\mathrm{S}}$, and the pressure flag $b=0,1$. We consider a fiducial model with $M=10^{8}\rm\,M_{\mathrm{\odot}}$$\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.1$, $l_{\mathrm{E}}=0.5$ and $b=0$. Of these parameters, Fig. 6 explores the effect of varying $\alpha$ and $l_{\rm E}$.
@@
We vary the Eddington ratio from $l_{\rm E}=10^{-3}$ to $l_{\rm E}=1$, capturing the range of observed AGNs (Suh et al., 2015). The Eddington ratio parameterizes the accretion rate, which plays a key role in the disc dynamics at all radial distances from the BH. Scaling relations in the optically thick regime far from the disc (see Sec. 2.1.4) indicate that the SG03 model maintains a constant temperature and density at $r\gtrsim 10^{5}R_{\mathrm{s}}$. Higher accretion rates leads to higher effective temperatures [Eq. (8)], higher disc temperatures overall [Eq. (12)], and higher total pressure in the disc [Eq. (9)], which also implies that $h$ must be higher to maintain hydrostatic equilibrium.
@@ 4 Disc migration
Migration in gas discs was first proposed based on formulae that considered how spiral-wave structures can be sustained in galaxies (Feldman & Lin, 1973; Goldreich & Tremaine, 1979; Lin & Papaloizou, 1979). An object orbiting in a gas disc exchanges angular momentum with its surroundings, leading to changes in its orbit and thus a net radial migration in the disc (Armitage, 2020). These migration torques were predicted by Goldreich & Tremaine (1979), described for planets in protoplanetary discs by Ward (1997), improved upon by Paardekooper & Mellema (2006); Paardekooper et al. (2010), studied for the case of planets by Nelson et al. (2000); Tanaka et al. (2002); Paardekooper & Mellema (2006); Kley & Crida (2008); Paardekooper et al. (2010); Lyra et al. (2010), and extended to the AGN case by Syer et al. (1991); Artymowicz et al. (1993); McKernan et al. (2011, 2012); Bellovary et al. (2016). A key phenomenon emerging from these studies is the potential occurrence of migration traps — locations in the gas disc where the net radial migration torque is zero. Depending on the mass ratio between the migrator, the central object or the disc, the migrator may clear a gap (referred to as Type II migration) or deplete material at the migration trap without clearing a gap (Type I migration) (Ward, 1997). In this work, we only consider the case of Type I migration. Migration traps are an effective mechanism for merging stellar-mass BH binaries embedded in AGN discs, especially in a hierarchical manner (McKernan et al., 2012; Bellovary et al., 2016; Yang et al., 2019; Santini et al., 2023; Vaccaro et al., 2023, see Gerosa & Fishbach 2021 for a review). Earlier works by Bellovary et al. (2016) and Grishin et al. (2023) showed that the the location of migration traps does not depend on the properties of the migrating object. The location of these migration traps turns out to be a non-trivial function of the AGN disc parameters, ultimately set by the complex interplay of the gradients of the surface density $\Sigma_{\mathrm{g}}$ and temperature $T$. Migration traps are thus an ideal context to showcase our implementation of the Sirko & Goodman (2003) and Thompson et al. (2005) disc models. Table 2 summarizes all parameters used for this section.
@@ 4.1 Torque implementation
In particular, we apply our AGN disc models to the methods by Grishin et al. (2023), adopting their migration torque and thermal torque expressions. Grishin et al. (2023) use a simpler AGN disc model where profiles are power laws in $M$, $r$ and accretion rate $\dot{M}$. Their discs are relatively similar to the SG03 models with $M=10^{6}M_{\odot}$ and $l_{\rm E}=0.1$. When using migration torques by Paardekooper et al. (2010) which assume the disc is locally isothermal, Grishin et al. (2023) report the existence of migration traps. However, migration traps disappear when considering the updated migration torque formulas by Jiménez & Masset (2017). Grishin et al. (2023) then add a new type of migratory torque, namely the thermal torque by Masset (2017), and find that migration traps are able to form in their AGN disc model once more. We apply the same methodology and formulas to our more complex AGN models.
@@
is the critical luminosity. If $L=L_{\mathrm{c}}$, the hot and cold torques acting on the migrator balance out and $\Gamma_{\mathrm{therm}}=0$. The size of the lobes $\lambda$ is given by (Grishin et al., 2023)
@@
and the corotation radius is (Grishin et al., 2023):
@@
Figure 8: The absolute values of the migration torques for a $m_{\rm BH}=10\,{\rm M}_{\odot}$ BH orbiting a $M=10^{6}M_{\odot}$ central BH in AGN discs. The left panels show torque profiles for a SG03 disc with $\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.1$, $l_{\mathrm{E}}=0.5$ and $b=0$. The right panels show torque profiles for a TQM05 disc with $\epsilon_{\mathrm{T}}=10^{-3}$, $\xi=1$, $m=0.2$, $r_{\rm out}=10^{7}R_{\mathrm{s}}$ and $\dot{M}_{\rm out}=1.5\times 10^{-2}\,M_{\odot}\,{\rm yr}^{-1}$. Migration torques, thermal torques, and their combination are shown in first three rows from the top, respectively. The bottom panel shows the midplane surface density of the disc for each case. For the Type I migration torques considered in the top row, we show both results using prescriptions by both Paardekooper et al. (2010) (light curves) and Jiménez & Masset (2017) (heavy curves). Colors indicate the sign of the torque, with blue referring to inward migration (i.e. positive torques) and orange referring to outward migration (i.e. negative torques). Vertical grey lines indicate the migration traps.

### v2

@@ 1 Introduction
Active galactic nuclei (AGNs) are compact regions at the center of galaxies powered by gas accretion onto supermassive black holes (BHs) as opposed to solely the radiation from stars. The underlying theory describing the accretion disc of AGNs was first introduced by Zel’dovich (1964) and Salpeter (1964). AGNs have been studied at low and high redshifts across several electromagnetic wavelengths, capturing a wide range of astrophysical phenomena (see Netzer 2015; Padovani et al. 2017; Hickox & Alexander 2018; Bianchi et al. 2022 for broad reviews on the topic). Due to the deep gravitational well surrounding the central BH, the gas in the accretion disc is expected to reach temperatures of ${\sim}10^{5}$ and surface densities of ${\sim}10^{5}$g cm-2. The accretion disc of the central BH extends to sub-pc scales and is surrounded by optically thick material which is coupled to the disc itself. These components are collectively referred to as the AGN disc, which is expected to extend to separations of 1-10 pc (Netzer, 2015). Because of high obscurations and uncertainty in observations, the actual size of AGN discs is somewhat unclear but tends to be larger than what is expected from theoretical models (Jha et al., 2022; Guo et al., 2022a, b).
@@
AGN discs are unique astrophysical environments with a rich phenomenology, including high-energy jets, dusty torii, and accreting BHs (Padovani et al., 2017). In the context of gravitational-wave observations, AGN discs are studied as host environments for compact-binary formation and mergers (McKernan et al., 2012, 2011; Yang et al., 2019; Secunda et al., 2019; Fabj et al., 2020; Tagawa et al., 2020; Trani et al., 2024). The large escape velocity around a supermassive BH implies that objects are likely to be retained in the disc vicinities, potentially forming a large population of stellar-mass BHs that have a higher likelihood of interacting. The dense gas in the disc can facilitate binary formation, accelerate the inspiral, and induce chains of hierarchical BH mergers (Gerosa & Fishbach, 2021; Santini et al., 2023; Whitehead et al., 2023; Vaccaro et al., 2023). The occurrence of hierarchical mergers in AGN discs crucially depends on the presence of the so-called migration traps, namely locations in the disc where the migration torque changes sign, which is still an open issue in AGN-disc modeling (Bellovary et al., 2016; Tagawa et al., 2020; Grishin et al., 2024).
@@
Early models of AGNs discs consist of one-dimensional, steady-state, semi-analytic solutions utilizing parametric prescriptions. Subsequent computational advancements allowed for models capturing more complex physics, such as radiative transfer, gas phase transitions, magnetic fields, and general relativity (e.g. Wilson, 1972; Hawley et al., 1984; Falle, 1991; Schartmann et al., 2005; Schartmann et al., 2008; Wada et al., 2009; Huško & Lacey, 2023). Nevertheless, one-dimensional models remain highly valuable today due to their computational efficiency and insightful perspectives on the structure of AGN discs. This makes them particularly useful in the study of interactions between compact objects and BHs. The first of these one-dimensional approaches dates back to Shakura & Sunyaev (1973), who first model geometrically thin, optically thick discs around a BHs. Building on this seminal work, two models are most commonly used in the field, namely those by Sirko & Goodman (2003) and Thompson et al. (2005) (but see also Goodman 2003; Levin 2003; Haiman et al. 2009; Cantiello et al. 2021; Gilbaum & Stone 2022; Grishin et al. 2024; Hopkins et al. 2024a, b). Both these models assume some heating mechanisms in the disc that marginally support the outer regions from collapsing due to self-gravity and formulate one-dimensional sets of equations for the AGN-disc profile as a function of a number of parameters such as the mass of the central BH and the accretion rate.
@@
The Sirko & Goodman (2003) and Thompson et al. (2005) models are widely used and underpin some of the key, qualitative results in the field of AGN-disc physics. Despite that, the underlying parameters and methods are often left unspecified. Achieving a stable numerical implementation of these disc solutions is not straightforward and codes in this area have not been released in the public domain. The goal of this paper is to critically re-analyze the AGN disc models by Sirko & Goodman (2003) and Thompson et al. (2005). In particular, we clarify the model equations one needs to solve (and crucially the order one needs to solve them), highlight the choices one has to make to obtain stable solutions, and provide a highly customizable implementation. Our software is made publicly available in a Python package called pAGN (short for “parametric AGNs”, pronounced as “pagan”).
@@ 2.1.2 Inner regime
The temperature profile in the disc depends on the optical depth $\tau_{\rm v}=\kappa\rho h$, where $\kappa(\rho,T)$ is the opacity. The latter is obtained using interpolated values by Semenov et al. (2003) when $T<10^{4}{\rm K}$ and Badnell et al. (2005) when $T>10^{4}{\rm K}$, the set of which we refer to as the “combined” opacity. These are newer prescriptions for the opacity compared to those by Iglesias & Rogers (1996) and Alexander & Ferguson (1994) used by Sirko & Goodman (2003). The opacities in Semenov et al. (2003) are calculated for silicate grains; the effect of graphite that is important at temperatures of ${\sim}2000$ K and may be responsible for the broad line region in AGN observations (see Baskin & Laor, 2018) is ignored. The inclusion of the effect of graphite in pAGN is left to future work. From the opacity and effective temperature, we look for solutions in $T$ by assuming the disc ring is in radiative equilibrium:
@@ 2.1.3 Outer regime
Figure 3: Radial profile in the SG03 disc model for the opacity $\kappa$, the mass density $\rho$, the pressure fraction $\beta$, the optical depth $\tau_{\rm v}$, the disc mass $M_{\rm disc}$, the sound speed $c_{\mathrm{s}}$, the temperature $T$, the most unstable mass $M_{\rm{most}\,\rm{unst.}}$, the Toomre stability parameter $Q_{\mathrm{S}}$, the effective temperature $T_{\rm eff}$, the surface mass density $\Sigma_{\mathrm{g}}$ and the half-thickness $h$. All input values are the same as in Fig. 2 by Sirko & Goodman (2003): $M=10^{8}M_{\odot}$, $\alpha=0.01$, $l_{\mathrm{E}}=0.5,\epsilon_{\rm S}=0.1$. Blue (orange) curves indicate the case where $b=0$ ($b=1$) and the viscosity is proportional to total (gas) pressure.
@@ 2.1.4 Disc profiles
In fig. 2 in Sirko & Goodman (2003), there are three different solutions for the disc parameters from $r\gtrsim 5\times 10^{5}R_{\rm s}$. Our implementation recovers the same behavior, but we only accept the continuous, high-temperature, low-opacity solution. For this case, the midplane temperature of the disc remains above $10^{3}$ K and the opacity drops to $10^{-3}$ cm2g-1 in the outer regime, both of which affect the gas and radiation pressure profiles, as reflected in the parameter $\beta$. The transition between the inner and outer disc regimes takes place at $r\approx 10^{3}R_{\mathrm{s}}$, which is consistent with the original results by Sirko & Goodman (2003).
@@ 2.2.4 Accretion criterion
Unlike Sirko & Goodman (2003), the Thompson et al. (2005) model presents an accretion rate $\dot{M}$ that changes as a function of the radial separation $r$. This naturally means that if the accretion rate at the outer boundary is too low, not enough gas is able to reach the central BH to maintain high temperatures and bright AGN luminosities, which are expected to be in the $10^{-3}-0.5\,L_{\rm Edd}$ range, see Heckman et al. (2004); Kollmeier et al. (2006); Suh et al. (2015); Kong & Ho (2018). This introduces a minimum threshold for $\dot{M}_{\rm out}$. Thompson et al. (2005) argue that accretion rates of ${\sim}1-10\rm\,M_{\odot}{\rm yr}^{-1}$ at the inner disc boundary $r_{\rm min}$ are sufficient to produce a bright AGN when the central BH mass is ${\sim}10^{9}\rm\,M_{\odot}$. This is equivalent to a minimum BH accretion rate of $\dot{M}\sim 0.2~{}\dot{M}_{\rm Edd}=0.2~{}L_{\rm Edd}/(0.1c^{2})$ at $r=r_{\rm min}$. Using Eq. (47) in Thompson et al. (2005), we find that over a wavelength range of $[10^{-8}\,\mathrm{m},10^{-3}\,\mathrm{m}]$, setting $0.2~{}\dot{M}_{\rm Edd}$ gives a disc bolometric luminosity of $2\times 10^{-4}\,L_{\rm Edd}$.
@@
We find that using $\dot{M}_{\mathrm{c}}$ from Eq. (26) as a threshold is too stringent and often omits signals that produce bright AGN discs. In the following, we use Eq. (27) as an initial guess for $\dot{M}_{\rm out}$ but then make adjustments if the accretion rate is not large enough to form a luminous AGN. Developing a full prescription is left to future work. As a precaution to avoid setting an $\dot{M}_{\rm out}$ that is too high, TQM05 suggests a maximum limit for $\dot{M}_{\rm out}$ equal to $\dot{M}_{\rm max}=8\pi\rho h\sigma^{2}r/\epsilon_{\rm T}c=L_{\rm M}/\epsilon_{\rm T}c^{2}$, where $L_{\rm M}$ is the limiting Eddington-like luminosity below which a galaxy will not have momentum driven winds that are high enough to significantly reduce the gas in the disc (Murray et al., 2005).
@@ 2.2.5 Disc profiles
Figure 4 reproduces Fig. 6 in Thompson et al. (2005), assuming either the Semenov opacities (as was done by Thompson et al. 2005) or the combined opacities. The input parameters are $\sigma=300$ km s-1, $M\approx 10^{9}\,M_{\odot}$ [instead of Eq. (15) we use the $M$-$\sigma$ relation by Tremaine et al. (2002) as was done by Thompson et al. (2005)], $\dot{M}_{\rm out}=320M_{\odot}$ yr-1, $r_{\rm out}=200$ pc, $m=0.2$, $\epsilon_{\mathrm{T}}=10^{-3}$, and $\xi=1$. Our results shown in Fig. 4 are generally in agreement with those by Thompson et al. (2005).
@@
Figure 4: Radial profile in TQM05 disc model for the temperature $T$, the effective temperature $T_{\rm eff}$, the mass density $\rho$, the surface mass density $\Sigma_{g}$, the gas pressure $p_{\rm gas}$, the radiation pressure $p_{\rm rad}$, the gas pressure fraction $\beta$, the half-thickness of the disc $h$, the sound speed $c_{\rm s}$, the opacity $\kappa$ and the optical depth $\tau_{\rm v}$. The input values have been chosen to reproduce Fig. 6 in Thompson et al. (2005): $\sigma=300$ km/s, $\epsilon_{\mathrm{T}}=10^{-3}$, $m=0.2$, $\dot{M}_{\rm out}=320M_{\odot}$ yr-1, and $r_{\rm out}=200$ pc. Models shown in blue use the opacities by Semenov et al. (2003), models shown in orange use the combined datasets from Semenov et al. (2003) and Badnell et al. (2005).
@@ 3.2 Input parameters
The SG03 model has five input parameters: the mass of the central BH $M$, the luminosity ratio $l_{\rm E}$ (or alternatively the accretion rate $\dot{M}$), the disc viscosity $\alpha$, the BH radiative efficiency $\epsilon_{\mathrm{S}}$, and the pressure flag $b=0,1$. We consider a fiducial model with $M=10^{8}\rm\,M_{\mathrm{\odot}}$$\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.01$, $l_{\mathrm{E}}=0.5$ and $b=0$. Of these parameters, Fig. 6 explores the effect of varying $\alpha$ and $l_{\rm E}$.
@@
We vary the Eddington ratio from $l_{\rm E}=10^{-3}$ to $l_{\rm E}=1$, capturing the range of observed AGNs (Heckman et al., 2004; Kollmeier et al., 2006; Suh et al., 2015; Kong & Ho, 2018). The Eddington ratio parameterizes the accretion rate, which plays a key role in the disc dynamics at all radial distances from the BH. Scaling relations in the optically thick regime far from the disc (see Sec. 2.1.4) indicate that the SG03 model maintains a constant temperature and density at $r\gtrsim 10^{5}R_{\mathrm{s}}$. Higher accretion rates leads to higher effective temperatures [Eq. (8)], higher disc temperatures overall [Eq. (12)], and higher total pressure in the disc [Eq. (9)], which also implies that $h$ must be higher to maintain hydrostatic equilibrium.
@@ 4 Disc migration
Migration in gas discs was first proposed based on formulae that considered how spiral-wave structures can be sustained in galaxies (Feldman & Lin, 1973; Goldreich & Tremaine, 1979; Lin & Papaloizou, 1979). An object orbiting in a gas disc exchanges angular momentum with its surroundings, leading to changes in its orbit and thus a net radial migration in the disc (Armitage, 2020). These migration torques were predicted by Goldreich & Tremaine (1979), described for planets in protoplanetary discs by Ward (1997), improved upon by Paardekooper & Mellema (2006); Paardekooper et al. (2010), studied for the case of planets by Nelson et al. (2000); Tanaka et al. (2002); Paardekooper & Mellema (2006); Kley & Crida (2008); Paardekooper et al. (2010); Lyra et al. (2010), and extended to the AGN case by Syer et al. (1991); Artymowicz et al. (1993); Levin (2003); McKernan et al. (2011, 2012); Bellovary et al. (2016); Derdzinski & Mayer (2023). A key phenomenon emerging from these studies is the potential occurrence of migration traps — locations in the gas disc where the net radial migration torque is zero. Depending on the mass ratio between the migrator, the central object or the disc, the migrator may clear a gap (referred to as Type II migration) or deplete material at the migration trap without clearing a gap (Type I migration) (Ward, 1997). In this work, we only consider the case of Type I migration. Migration traps are an effective mechanism for merging stellar-mass BH binaries embedded in AGN discs, especially in a hierarchical manner (McKernan et al., 2012; Bellovary et al., 2016; Yang et al., 2019; Santini et al., 2023; Vaccaro et al., 2023, see Gerosa & Fishbach 2021 for a review). Earlier works by Bellovary et al. (2016) and Grishin et al. (2024) showed that the the location of migration traps does not depend on the properties of the migrating object. The location of these migration traps turns out to be a non-trivial function of the AGN disc parameters, ultimately set by the complex interplay of the gradients of the surface density $\Sigma_{\mathrm{g}}$ and temperature $T$. Migration traps are thus an ideal context to showcase our implementation of the Sirko & Goodman (2003) and Thompson et al. (2005) disc models. Table 2 summarizes all parameters used for this section.
@@ 4.1 Torque implementation
In particular, we apply our AGN disc models to the methods by Grishin et al. (2024), adopting their migration torque and thermal torque expressions. Grishin et al. (2024) use a simpler AGN disc model where profiles are power laws in $M$, $r$ and accretion rate $\dot{M}$. Their discs are relatively similar to the SG03 models with $M=10^{6}M_{\odot}$ and $\alpha=0.01$. When using migration torques by Paardekooper et al. (2010) which assume the disc is locally isothermal, Grishin et al. (2024) report the existence of migration traps. However, migration traps disappear when considering the updated migration torque formulas by Jiménez & Masset (2017). Grishin et al. (2024) then add a new type of migratory torque, namely the thermal torque by Masset (2017), and find that migration traps are able to form in their AGN disc model once more. We apply the same methodology and formulas to our more complex AGN models.
@@
is the critical luminosity. If $L=L_{\mathrm{c}}$, the hot and cold torques acting on the migrator balance out and $\Gamma_{\mathrm{therm}}=0$. We approximate the luminosity of the migrator, $L$, to be its Eddington luminosity [see Eq. (3), replacing the mass $M$ with the mass of the migrator $m_{\rm BH}$]. The size of the lobes $\lambda$ is given by (Grishin et al., 2024)
@@
and the corotation radius is (Grishin et al., 2024):
@@
The thermal torque given by Eq. (34) is expected to diminish in optically thin discs. Following Grishin et al. (2024), we multiply Eq. (34) by a factor of $1-\exp{-\lambda\tau_{\rm v}/h}$ ). Additionally, when the mass of the migrator exceeds the thermal mass $m_{\rm th}$, the thermal torque will be reduced (Guilera et al., 2021). The thermal mass is defined by:
$\frac{m_{\rm th}}{m_{\rm BH}}=\frac{\chi}{c_{\rm s}R_{\rm B}}\,,$ (38)
where $R_{\rm B}$ is half the Bondi radius
$R_{\rm B}=\frac{Gm_{\rm BH}}{c_{\rm s}^{2}}\,.$ (39)
In the regions where $h<R_{\rm B}$ we use the disk height $h$ in place of half the Bondi radius $R_{\rm B}$. To correct for the critical thermal mass, we split Eq. 34 into its heating component (the positive $L/L_{\rm c}$ term) and its cooling component (the negative term), which we label as $\Gamma_{\rm therm,\,hot}$ and $\Gamma_{\rm therm,\,cold}$ respectively. The total thermal torque is described by Eq. 34 unless $\mu_{\rm th}\equiv m_{\rm th}/m_{\rm BH}<1$. In regions of the disc where $\mu_{\rm th}<1$, the thermal torque is instead given by:
$\Gamma_{\rm therm}=\Gamma_{\rm therm,\,hot}\frac{4\mu_{\rm th}}{1+4\mu_{\rm th}}+\Gamma_{\rm therm,\,cold}\frac{2\mu_{\rm th}}{1+2\mu_{\rm th}}$ (40)
which is an approximation of numerical fits detailed in Velasco Romero & Masset (2020) and used in Guilera et al. (2021); Grishin et al. (2024).
Figure 8: The absolute values of the migration torques for a $m_{\rm BH}=10\,{\rm M}_{\odot}$ BH orbiting a $M=10^{6}M_{\odot}$ central BH in AGN discs. The left panels show torque profiles for a SG03 disc with $\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.01$, $l_{\mathrm{E}}=0.5$ and $b=0$. The right panels show torque profiles for a TQM05 disc with $\epsilon_{\mathrm{T}}=10^{-3}$, $\xi=1$, $m=0.2$, $r_{\rm out}=10^{7}R_{\mathrm{s}}$ and $\dot{M}_{\rm out}=1.5\times 10^{-2}\,M_{\odot}\,{\rm yr}^{-1}$. Migration torques, thermal torques, and their combination are shown in first three rows from the top, respectively. The bottom panel shows the midplane surface density of the disc for each case. For the Type I migration torques considered in the top row, we show both results using prescriptions by both Paardekooper et al. (2010) (light curves) and Jiménez & Masset (2017) (heavy curves). Colors indicate the sign of the torque, with blue referring to inward migration (i.e. positive torques) and orange referring to outward migration (i.e. negative torques). Vertical grey lines indicate the migration traps for all torque prescriptions except for Paardekooper et al. (2010) in the top panel.